OpenStudy (anonymous):

A large diamond with a mass of 712.04 grams was recently discovered in a mine. If the density of the diamond is 3.51 grams over centimeters cubed, what is the volume? Round your answer to the nearest hundredth. 202.9 cm3 507.02 cm3 717.06 cm3 3574.44 cm3

OpenStudy (baru):

\(density=\frac{mass}{volume}\)

OpenStudy (anonymous):

So would you do this: \[density = \frac{ x}{ 712.04 }\] ? I don't understand if it says that denisty is already found (3.51)...?

OpenStudy (anonymous):

Oh wait \[3.51 = \frac{ x }{ 712.04 }\] ?

OpenStudy (baru):

the idea you have is correct, but you have mixed up mass and volume. x should be in the denominator and 712.04 in the numerator

OpenStudy (anonymous):

So would it be: \[3.15 = \frac{ 712.04 }{ x }\] But how would you solve that..?

OpenStudy (anonymous):

*oops 3.51

OpenStudy (baru):

you used 'x' to represent volume. so you have been thought some basic algebra?

OpenStudy (anonymous):

Yes. Algebra 1, this is Geometry though.. Lol.

OpenStudy (baru):

you have to multiply both sides of the equation by 'x'

OpenStudy (anonymous):

So you'd multiply 'x' by 712.04? Wouldn't you get the same answer (712.04) I'm a little lost.. Or would you multiply the 3.51 to 712.04?

OpenStudy (baru):

\[x \times 3.15 = \frac{x\times 712.04}{x}\]

OpenStudy (anonymous):

Yes but if you have a 'x', doesn't it just equal to 1? So '1x' ?

OpenStudy (baru):

i didnt understand what you mean. according to you, what does 'x' represent?

OpenStudy (anonymous):

Well, doesn't 1 represents x?

OpenStudy (baru):

No

OpenStudy (anonymous):

So how would you find the 'x' in this problem then?

OpenStudy (baru):

\(density=\frac{mass}{volume}\) we know density, we know mass... but we do not know volume . so we have used the letter 'x' like a shortform for 'volume' and substituted into the above formula. so 'x' represents 'volume' and volume is the quantity that the question is asking us to find

OpenStudy (baru):

u understand that?

OpenStudy (anonymous):

Yes, I understand that. But isn't the volume going to be one of the answer choices? ^

OpenStudy (baru):

all the options given are possible values for 'volume' but only one of them is correct. we have to find which one by using our equation

OpenStudy (anonymous):

So, I'm guessing you would take each 'volume' and replace it for 'x' to see which one is right in the equation?

OpenStudy (baru):

yes that's definitely one way to do it :)

OpenStudy (baru):

have you learnt how to solve algebraic equations?

OpenStudy (unklerhaukus):

density = mass / volume we can rearrange the equation by multiplying both side by the volume and, density * volume = (mass / volume) * volume = mass dividing both side by the density density * volume / density = mass / density we get volume = mass / density

OpenStudy (anonymous):

Well yeah, but I'm kinda forgetful.. I know how to solve most but this one is giving me problems xD. So would the next step be this: \[202.9 \times 3.51 = \frac{ 202.9 \times 712.04 }{ x }\] ?

OpenStudy (unklerhaukus):

don't put the numbers in till the equation is solved

OpenStudy (anonymous):

So you would divide everything by 3.51?

OpenStudy (unklerhaukus):

We start the definition of density density = mass / volume we rearrange the equation so solve to volume (as the question asks for) volume = mass / density now we use the values mass = 712.04 g density = 3.51 g/cc plug these values into the equation

OpenStudy (anonymous):

So, \[x = \frac{ 712.04 }{ 3.51 } ?\]

OpenStudy (unklerhaukus):

yeahs (but don't forget the units )

OpenStudy (anonymous):

So would you start it of by multiplying or dividing? Or is that not the next step.. ?

OpenStudy (unklerhaukus):

compute (712.04 g) ÷ (3.51 g/cc)

OpenStudy (anonymous):

Okay let me do that, give me a minute :). Thanks for helping me

OpenStudy (anonymous):

I got 202.86 making sure its g and g/cc

OpenStudy (unklerhaukus):

the units g / (g/cc) = 1/(1/cc) = cc = cubic centimetres

OpenStudy (unklerhaukus):

the grams cancel away

OpenStudy (anonymous):

So, would it be the 202? @UnkleRhaukus

OpenStudy (unklerhaukus):

if you round to the nearest hundredth, you answer won't match the options

OpenStudy (anonymous):

So you would round that 202.8? @UnkleRhaukus

OpenStudy (unklerhaukus):

i would have had 202.86 cm^3,

OpenStudy (anonymous):

So since it's cm^3 would you multiply the 202.86^3? I think that's wrong..

OpenStudy (unklerhaukus):

(the ^3 is only on the unit)

OpenStudy (unklerhaukus):

one could argue that the last two digits are invalid, as 3.51 has only three significant figures of accuracy , so that 203 cm^3, is more true

OpenStudy (anonymous):

This is confusing, what would be your next step then to solving the equation?

OpenStudy (unklerhaukus):

it's multiple choice right, one of the answers is correct, we followed the right steps and got an answer that was almost exactly a match, (all the other options are way off)

OpenStudy (anonymous):

So I was right about the 202.8? We just had to round it up, and it should be close to the first option?

OpenStudy (unklerhaukus):

You were right with 202.86 cm3 Best match is A 202.9 cm3 (we have to ignore the rounding error i guess )

OpenStudy (anonymous):

Okay thank you for helping me solve the problem! Sorry it took so long for me to figure out :)

OpenStudy (anonymous):

We were right by the way :) The answer was correct. @UnkleRhaukus

OpenStudy (unklerhaukus):

do i get to keep the diamond?

OpenStudy (anonymous):

Aha, sure :P. @UnkleRhaukus